OpenStudy (anonymous):

If sinθ=-1/4 and θ terminates in third quad., find the exact value of sin 2θ

OpenStudy (anonymous):

???

OpenStudy (anonymous):

Okie, found it :)

OpenStudy (anonymous):

\[\sin²(\theta)=1/16=[1-\cos(2.\theta)]/2\] \[\cos(2.\theta)=1-1/8\] \[\sin(2.\theta)=-\sqrt{1-\cos(2.\theta)^{2}}\]

OpenStudy (anonymous):

sin2x = 2sinxcosx cos^2x = 1 - sin^2x = 1 - ( -1/4)^2

OpenStudy (anonymous):

Can you follow?

OpenStudy (anonymous):

the "-" is due to third quad

OpenStudy (anonymous):

@victor do you need more help?

OpenStudy (anonymous):

i dont understand how you did that?

OpenStudy (anonymous):

i understand sin2x = 2sinxcosx but then how did you get cos^2x = 1 - sin^2x = 1 - ( -1/4)^2 @Chlorophyll

OpenStudy (anonymous):

Given: sinx, and theta in 3rd quadrant Ask : sin2x = 2sinxcosx

OpenStudy (anonymous):

=> need cosx

OpenStudy (anonymous):

Sin^2x + cos^2x = 1

OpenStudy (anonymous):

ohhh i see it. so youre not using the double angle you just replaced it

OpenStudy (anonymous):

cosx = sqrt ( 1 - sin^2x ) I don't want to confuse you, so I go step by step!

OpenStudy (anonymous):

Are you okie up to this step?

OpenStudy (anonymous):

yea cosx=1+sin(1/2)

OpenStudy (anonymous):

NO, that's different formula, half angle!

OpenStudy (anonymous):

cosx = sqrt ( 1 - sin^2x ) You want angle is x not x/2, not 2x

OpenStudy (anonymous):

yea but isnt x -1/4 so sqrt 1/4 is 1/2

OpenStudy (anonymous):

cosx = sqrt ( 1 - sin^2x ) = sqrt ( 1 - 1/16 ) = sqrt ( 15/16) = sqrt (15) / 4

OpenStudy (anonymous):

oh i sqrt first. you cant just leave it as sqrt 15/16

OpenStudy (anonymous):

and then you plug in your sqrt to cos

OpenStudy (anonymous):

That's how you find cosx from sinx ( very traditional way)

OpenStudy (anonymous):

so whats the final answer?

OpenStudy (anonymous):

Notice that x belongs the THIRD quadrant -> both cosx and sinx negative => cosx = - sqrt (15)/4 Now plug it into sin2x = 2 sinx * cosx = ....

OpenStudy (anonymous):

I assume you're fine here, all you do is multiply 2 sinx * cosx

OpenStudy (anonymous):

\[\sqrt{15/4} or \sqrt{15/16}\]

OpenStudy (anonymous):

for cos

OpenStudy (anonymous):

\[cosx = \sqrt{15}/ 4\]

OpenStudy (anonymous):

\[\sin (2 \theta )=2 \sin (\theta ) \cos (\theta )=2 \sin (\theta ) \sqrt{1-\sin ^2(\theta )}=\frac{2}{4} (-1) \sqrt{1-\frac{1}{16}}=-\frac{\sqrt{15}}{ 8} \]